OpenStudy (anonymous):

a triangle is formed by the x axis the y axis and the line y=ax+b if a=-b then what is the volume of the cone generated by rotating this triangle around y axis.

OpenStudy (anonymous):

Figure out how high the cone is. That's the y-intercept, where x = 0. (Set x = 0 and solve for y). Figure out the radius of the cone's base. That's the (absolute value of the) x-intercept, where y = 0. Volume of a cone is (1/3)pi*r^2*h

OpenStudy (anonymous):

what is the answer? @CO_oLBoY

OpenStudy (anonymous):

\[\frac{ \pi }{ 3 } a\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

y=ax+b and a=-b or b=-a so, y=ax-a y=a(x-1)

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

@CO_oLBoY

OpenStudy (anonymous):

y = 0 then x = 1 = r V = (1/3)*pi*1*a = (pi /3) a

OpenStudy (anonymous):

ok but why x=r?

OpenStudy (anonymous):

coz it was rotating around y-axis hence making radius x

OpenStudy (anonymous):

and why h=a?

OpenStudy (anonymous):

coz its on y

OpenStudy (anonymous):

mmm ok ty @CO_oLBoY

OpenStudy (anonymous):

yw